OpenStudy (anonymous):

Determine the value of k so that the lines kx-y=-1/3 and 3y=1-6x will intersect at 1 point?

OpenStudy (ikram002p):

:) solve the system of equations

OpenStudy (anonymous):

Can't we go by equating coefficients only?

OpenStudy (ikram002p):

3y+6x-1=0 -y+kx+1/3=0

OpenStudy (anonymous):

how do that this equation?

OpenStudy (ikram002p):

thus , add first equation to 3* second equation 3y+6x-1=0 -3y+3kx+1=0 so 3k=6

OpenStudy (ikram002p):

typo -3k=6

OpenStudy (anonymous):

then???

OpenStudy (anonymous):

Try to find out k there by dividing by -3..

OpenStudy (anonymous):

k= -2

OpenStudy (anonymous):

Yep.. :)

OpenStudy (anonymous):

k= -2 this a final answer

OpenStudy (anonymous):

According to Ikram, Yes this is true.. :)

OpenStudy (anonymous):

Determine the value of k so that the lines kx-y=-1/3 and 3y=1-6x will not intersect?

OpenStudy (anonymous):

Is this other question or the same question?

OpenStudy (anonymous):

other question